OpenStudy (anonymous):

Can someone help explain this to me? (hold on i need to write it up) thanks!

OpenStudy (anonymous):

\[Complete the proportions: 1. KM/KL= ?/JK 2. JM/?= JK/JL 3. ?/LK=Lk/LM 4. JM/?=KM/LM 5. LK/LM = JK/? 6. ?/JK=MK/MJ\]

OpenStudy (anonymous):

oh no it didn't come out right...ugh

OpenStudy (anonymous):